1) A research study wanted to examine whether religious belief and gender are related in their...

1)

  1. A research study wanted to examine whether religious belief and gender are related in their state. They selected a representative sample of men and women in their State.

    A. Select the null and alternative hypothesis of the test: ____ (enter option I or II)

    I) H0: there is a relationship between gender and belief about human origins
    Ha: there is no relationship between gender and belief about human origins
    II) H0: there is no relationship between gender and belief about human origins
    Ha: there is a relationship between gender and belief about human origins

    B. Results of the study are summarized in this table:

    Belief about human origins
    Religious texts Evolutionary Theory Total
    Gender Male 30 36 66
    Female 106 89 195
    Total 136 125 261

    The degrees of freedom for the the Chi-square statistic are: ____ (enter a number: 1,2,3 or 4)

    The test statistic is X-squared =____, with a p-value = ____. (round answers to the nearest hundredth. Do not enter the 0 before the decimal point.)

    C. The conclusion of the test is: ____ (enter an option: a, b, c or d from the list below)

    1. Gender and belief about human origins are independent.
    2. Gender and belief about human origins are related.
    3. There is enough statistical evidence to suggest that gender and belief about human origins are related.
    4. There is not enough statistical evidence to suggest that gender and belief about human origins are related.

2)

  1. What conditions need to be met to run a Chi-square test of independence?

    a.

    Variables must be related

    b.

    Variables must be independent

    c.

    Cell counts must be at least 5

    d.

    Expected cell counts must be at least 5
Answer #1

1A)

H0: there is no relationship between gender and belief about human origins

Ha: there is a relationship between gender and belief about human origin

b)

degree of freedom(df) =(rows-1)*(columns-1)= 1

applying chi square test:

Observed Oi religious evolutionary Total
male 30 36 66
female 106 89 195
total 136 125 261
Expected Ei=row total*column total/grand total religious evolutionary Total
male 34.39 31.61 66
female 101.61 93.39 195
total 136 125 261
chi square    χ2 =(Oi-Ei)2/Ei religious evolutionary Total
male 0.5606 0.6099 1.171
female 0.1897 0.2064 0.396
total 0.750 0.816 1.567

test statistic X2 =1.57 and p value =0.21

c)

There is not enough statistical evidence to suggest that gender and belief about human origins are related.

2) Expected cell counts must be at least 5